# DIVER-1: Deep Integration of Vast Electrophysiological Recordings at Scale

### v1

@@ Abstract
Electrophysiology signals such as EEG and iEEG are central to neuroscience, brain–computer interfaces, and clinical applications, yet existing foundation models remain limited in scale despite clear evidence that scaling improves performance. We introduce DIVER-1, a family of EEG and iEEG foundation models trained on the largest and most diverse corpus to date—5.3k hours of iEEG and 54k hours of EEG (1.6M channel-hours from over 17.7k subjects)—and scaled up to 1.82B parameters. We present the first systematic scaling law analysis for this domain, showing that they follow data-constrained scaling laws: for a given amount of data and compute, smaller models trained for extended epochs consistently outperform larger models trained briefly. This behavior contrasts with prior electrophysiology foundation models that emphasized model size over training duration. To achieve strong performance, we also design architectural innovations including any-variate attention, sliding temporal conditional positional encoding, and multi-domain reconstruction. DIVER-1 iEEG and EEG models each achieve state-of-the-art performance on their respective benchmarks, establishing a concrete guidelines for efficient scaling and resource allocation in electrophysiology foundation model development. [^1] Code available at: https://anonymous.4open.science/r/DIVER-1
@@ 1 Introduction
Scaling has been a fundamental driver of progress in artificial intelligence, from early perceptrons to modern large language models. Systematic increases in data, compute, and model size have yielded reliable performance gains across language and vision, motivating principled investigations of neural scaling laws (kaplan2020scaling; vision_scaling_zhai2022scaling).
Electrophysiology (Ephys)–including intracranial (iEEG) and scalp EEG–presents a distinct opportunity for foundation modeling in neuroscience, BCI, and clinical applications. The domain is marked by heterogeneity across subjects, recording sessions, montages, and neural states (eeg_hetereogenity_ebadi2025beyond). Recent EEG foundation models (EFMs) [^2] Although other terms such as large brain models (LBM) exist, we use EFM as it explicitly ties our setting to EEG (both iEEG and scalp EEG) and emphasizes that it is a foundation, not necessarily a large, model. leverage self-supervised pretraining over large unlabeled corpora and have reported consistent improvements on downstream decoding tasks (jiang2024labram; zhang2023brant; wang2024cbramod).
Despite rapid progress, two gaps remain. First, no research has scaled EFMs beyond the levels necessary to fully exploit this potential—previous efforts have been limited by computational resources and dataset availability. Second, no systematic, quantitative analysis of EFM scaling behavior has been conducted, leaving fundamental questions unanswered about optimal resource allocation and scaling strategies for neural data—in particular, how to allocate a fixed compute budget across model size and number of training epochs.
We address these gaps of limited scale exploration and the lack of quantitative scaling analysis by conducting the first systematic investigation of scaling laws in EFMs while pushing the boundaries of scale across all dimensions. Specifically, we expand across: (1) Data—assembling pretraining corpora with more than 77× more than the previous iEEG state-of-the-art model chau2025popt and 1.2× the channel-hours of iEEG data compared to zhang2023brant, plus about 10× more EEG data than existing state-of-the-art (wang2024cbramod); (2) Compute—utilizing substantially more computational resources than previous EFM studies; and (3) Model size—systematically evaluating architectures ranging from 13M to 1.82B parameters, 3× the size of the largest open-source iEEG foundation model (zhang2023brant)
Through this comprehensive scaling effort, we make a key discovery: EFMs precisely follow the data-constrained scaling laws of data_constrained_scaling_muennighoff2023scaling. Unlike classical Kaplan-style scaling (kaplan2020scaling; hoffmann2022chinchilla)—which assumes unlimited data and thus optimizes for one-epoch training—Ephys is inherently data-limited and requires multi-epoch training. The data-constrained framework of data_constrained_scaling_muennighoff2023scaling generalizes Kaplan-style scaling to this setting, and our empirical scaling curves quantitatively align with its predicted isoFLOPs trade-offs and exponents. This constraint fundamentally alters the optimal scaling strategy—we find that for a fixed data and compute budget, large models (≥1B parameters) trained for only a few epochs underperform smaller models trained for extended epochs, as additional training passes enable more effective utilization of the limited data.
Building on these insights, our experiments yield DIVER-1, a family of EEG and iEEG EFMs that achieves state-of-the-art performance across diverse neural decoding tasks. To maximize performance, the models incorporate architectural adaptations tailored for Ephys signals. The model employs sliding temporal conditional positional encoding for context-aware positioning while preserving channel-permutation equivariance, and any-variate attention mechanisms to handle variable electrode configurations with full spatio-temporal awareness. Additionally, spatiotemporal register tokens provide dedicated computational space without interfering with signal representations, while multi-domain reconstruction heads enable robust learning across temporal and spectral views without altering the encoder backbone.
The contributions of this work are fourfold:
  • DIVER-1 model family: We introduce a family of Ephys foundation models (EFMs) for both iEEG and EEG that achieve state-of-the-art performance in their respective modalities by leveraging our scaling insights and architectural innovations.
  • First systematic scaling law analysis for EFMs: We provide the first quantitative characterization of how EFMs scale with data, compute, and model size, revealing that EFMs follow data-constrained scaling laws (data_constrained_scaling_muennighoff2023scaling) due to the inherent scarcity of Ephys data. Building on this analysis, we show that under a fixed data and compute budget, smaller models trained for more epochs consistently outperform larger models trained briefly—offering clear guidance on how to allocate compute between model size and training epochs for efficient EFM development.
  • Unprecedented scale demonstration: We scale EFM pretraining to previously unattained levels across data volume, model size, and compute.
  • Novel architectural innovations for Ephys: We develop specialized components including any-variate attention mechanisms, sliding temporal conditional positional encoding (STCPE), register tokens, and multihead prediction architectures—enabling effective scaling analysis and performance gains.
Importantly, our findings show that prior approaches emphasizing model size as the primary axis of scaling are not well aligned with the realities of Ephys data (Appendix G). A more effective path under limited compute budgets is to prioritize training duration and subject diversity, a perspective that reframes how future EFMs should be developed.
2 Architecture
DIVER-1 uses an architecture custom-designed for multimodal EEG data that enables effective self-supervised pretraining through masked patch reconstruction, as described in Figure 1. This architecture consists of four main components: (1) patch encoding, (2) embedding enhancement, (3) a stack of MOIRAI blocks (woo2024moirai), and (4) multi-output projection. During pretraining, we randomly mask 50% of the input patches. The model then reconstructs missing information across multiple signal domains through the projection layer. The details of pretraining and architectural hyperparameters can be found in Appendix B.5.
@@ 2 Architecture
2.1 Structure Overview
@@
Patch encoding converts the raw signal into patch representations and extracts temporal features. Given an input time series $\mathbf{Y}\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels (variables) and $T$ is the sequence length, we patchify $\mathbf{Y}$ into 1 or 0.1 second patches, each containing $P\in\{500,50\}$ timepoints (at 500Hz sampling rate). During pretraining, $50\%$ of the patches are randomly masked with zero, and the remaining patches are passed through a three layer patch-wise convolutional neural network (CNN) to extract temporal features. This results in a basic representation $\mathbf{Y}_{\text{CNN}}\in\mathbb{R}^{C\times N\times d_{\text{model}}}$, where $N=T/P$ is the number of patches and $d_{\text{model}}$ is the embedding dimension.
Embedding enhancement augments the patch representations with signal and channel information through two mechanisms. First, spectral embeddings inject patch-wise frequency domain information, and channel position and modality embeddings jointly encode spatial electrode locations and electrode characteristics. Second, spatio-temporal conditional positional embedding (STCPE) is applied via a sliding window mechanism to capture both spatial channel relationships and local temporal information across the electrode array. The final transformer input is computed as:
MOIRAI blocks form the computational backbone of DIVER-1, adapted from MOIRAI (woo2024moirai) to model spatio-temporal dependencies in Ephys data. Our implementation incorporates three components: (1) Any-Variate Attention for improved multivariate attention, (2) Gated Linear Units (GLU) (shazeer2020glu) for improved expressivity through gating mechanisms, and (3) Spatio-Temporal Register Tokens (our novel addition) for dedicated computational space without interfering with EEG signal representations, inspired by vision_registers_darcet2023vision. The enhanced embedding $\mathbf{X}\in\mathbb{R}^{C\times N\times d_{\text{model}}}$ is processed through stacked MOIRAI blocks that preserve input dimensions while progressively refining spatio-temporal representations.
Multi-output projection enables simultaneous reconstruction across multiple signal domains during pretraining. Refined spatio-temporal representations are linearly projected to reconstruct patches in three complementary domains: raw time series, FFT, and STFT. This multi-domain reconstruction objective encourages comprehensive learning of temporal and spectral properties.
2.2 Key Component Details
Embedding enhancement components
@@ 2.2 Key Component Details
Spectral embeddings $\mathbf{E}_{\text{spectral}}$ inject patch-wise frequency information by applying FFT to each patch and projecting the features to $d_{\text{model}}$ dimensions via a learnable linear layer.
@@
Channel position and modality embeddings jointly encode spatial electrode locations and electrode modality to handle heterogeneous recording setups. While EEG follows standardized montages, iEEG channels are implanted in specific brain regions based on individual clinical needs. To address this heterogeneity, we encode both channel position embeddings $\mathbf{E}_{\text{position}}$ using the 3D spatial coordinates (registered to MNI space) and modality embeddings $\mathbf{E}_{\text{modality}}$ to distinguish between different electrode types. We then define the combined position–modality embedding term as:
$\mathbf{E}_{\text{position}},\mathbf{E}_{\text{modality}}]=\big[\,PE^{(x)}(x),\;PE^{(y)}(y),\;PE^{(z)}(z),\;\mathbf{E}_{\text{modality}}^{\text{type}}+\mathbf{E}_{\text{modality}}^{\text{subtype}}\,\big]$ (2)
where electrode coordinates $(x,y,z)$ are in centimeters and:
@@
for $j\in\{x,y,z\}$ and $d_{x}=d_{y}=d_{z}=d_{\text{model}}/4$, following PopT (chau2025popt). If the coordinates x,y,z for a given channel are not available, the positional embedding for that channel is set to zero. The modality embedding $\mathbf{E}_{\text{modality}}$ is computed as $\mathbf{E}_{\text{modality}}^{\text{type}}+\mathbf{E}_{\text{modality}}^{\text{subtype}}$, where $\mathbf{E}_{\text{modality}}^{\text{type}}$ distinguishes between EEG and iEEG channels, and $\mathbf{E}_{\text{modality}}^{\text{subtype}}$ encodes electrode subtypes (grid, strip, or depth for iEEG electrodes) through learnable embedding vectors shared across all patches within the same category.
Spatio-Temporal Conditional Positional Embedding (STCPE) addresses the need for dynamic positional encodings that can adapt to heterogeneous electrode configurations. Transformers require positional information because self-attention is inherently permutation-invariant. While vision transformers typically rely on fixed positional encodings, Conditional Positional Encoding (CPE) (chu2021cpe) applied lightweight convolutions over local neighborhoods, allowing positional information to adapt to the structure present in each input.
Building on this encoding scheme, Asymmetric Conditional Positional Encoding (ACPE) (wang2024cbramod) applied asymmetric convolutions across channels and time for EEG, but these convolutions operate along fixed channel axes and hence does not maintain channel permutation equivariance—the learned spatial relationships depend on the training-time channel order. This is a critical limitation for Ephys, where models must be invariant to arbitrary electrode reorderings to generalize across heterogeneous montages, recording systems, and channel counts.
@@
$\mathbf{E}_{\text{STCPE}}[:,t,:]=P_{\uparrow}\!\left(\sum_{k=-m}^{m}\mathbf{H}_{t+k}[:,\,k+m,:]\right)\in\mathbb{R}^{C\times w\times(d_{\text{model}})},\qquad t=1,\dots,N.$ (4)
@@
MOIRAI block components
Any-variate attention (woo2024moirai) enables adaptive spatial modeling across heterogeneous electrode configurations while maintaining full spatio-temporal attention capabilities. Unlike vanilla attention mechanisms that rely solely on input embeddings to differentiate between tokens and model their relationships, any-variate attention directly embeds spatio-temporal information into the attention computation itself through two main components: Rotary Position Embedding (RoPE) (su2024roformer) and binary attention bias.
@@
$E_{ij,mn}=(\mathbf{W}^{Q}\bm{x}_{i,m})^{T}\mathbf{R}_{i-j}(\mathbf{W}^{K}\bm{x}_{j,n})+u^{1}\cdot\mathbbm{1}_{\{m=n\}}+u^{2}\cdot\mathbbm{1}_{\{m\neq n\}}$ (5)
@@
Spatio-temporal register tokens are inspired by register tokens in vision transformers vision_registers_darcet2023vision. They consist of three types of learnable tokens each for channel, temporal, and combined spatio-temporal information, transforming input shape from $C\times N\times d_{\text{model}}$ to $(C+1)\times(N+1)\times d_{\text{model}}$. This transformation provides dedicated computational space to perform auxiliary computations without corrupting the primary Ephys signal representations.
@@ 3.1 Pretraining
DIVER-1 employs self-supervised pretraining based on masked patch reconstruction to learn robust representations of Ephys signals.
Multi-domain reconstruction objective. Rather than reconstructing only raw time series, DIVER-1 utilizes a multi-output projection architecture that maps each learned patch representation $\mathbf{h}_{c,n}\in\mathbb{R}^{d_{\text{model}}}$ to three complementary signal domains through parallel linear transformations. For each patch $(c,n)$, the representation is projected to: (1) raw time series $\hat{\mathbf{y}}_{c,n}^{\text{raw}}\in\mathbb{R}^{P}$ to reconstruct temporal dynamics, (2) FFT coefficients $\hat{\mathbf{y}}_{c,n}^{\text{FFT}}\in\mathbb{R}^{P/2+1}$ to capture frequency domain characteristics, and (3) STFT spectrogram $\hat{\mathbf{y}}_{c,n}^{\text{STFT}}\in\mathbb{R}^{F\times T_{s}}$ to model time-frequency relationships, with specific FFT and STFT parameters detailed in the implementation section in Appendix B.5.
@@
$\mathcal{L}_{\text{total}}=\sum_{(c,n)\in\mathcal{M}}\left[\lambda_{1}\mathcal{L}_{\text{MSE}}(\mathbf{y}_{c,n}^{\text{raw}},\hat{\mathbf{y}}_{c,n}^{\text{raw}})+\lambda_{2}\mathcal{L}_{\text{MSE}}(\mathbf{y}_{c,n}^{\text{FFT}},\hat{\mathbf{y}}_{c,n}^{\text{FFT}})+\lambda_{3}\mathcal{L}_{\text{MSE}}(\mathbf{y}_{c,n}^{\text{STFT}},\hat{\mathbf{y}}_{c,n}^{\text{STFT}})\right]$ (6)
where $\mathcal{M}$ denotes the set of masked patch indices. We used $(\lambda_{1,2,3})=(1,0.1,1)$ for $P=500$ (1 s patches) and $(1,1,0)$ for $P=50$ (0.1 s patches). These coefficients were chosen so that the different reconstruction losses operated on comparable numerical scales; for $P=50$, the window is too short for a meaningful STFT, so only the FFT term was used. This multi-domain approach encourages the model to learn comprehensive representations that capture both temporal dynamics and spectral features essential for robust Ephys understanding.
@@
Pretraining dataset. DIVER was pretrained on, to our knowledge, the largest and most diverse collection of Ephys datasets compared to previous EFMs. Our pretraining data encompasses diverse recording conditions including task-based experiments, resting-state recordings, and sleep studies, ensuring robust representation learning across different brain states. The pretraining datasets are summarized in Table 2 with further details in Appendix F.1.
All data underwent QAQC, minimal preprocessing, and resampling with the goal of preserving as much of the original signal as possible. QAQC applied conservative amplitude clipping, removing electrodes only when $3.33\%$ of samples exceeded the clipping threshold and discarding whole segments only when $50\%$ of channels were affected, ensuring minimal data loss while preventing extreme values from destabilizing training. Preprocessing then normalized EEG and iEEG amplitudes (100 µV and 200 µV scales, respectively), applied minimal filtering (0.3–0.5 Hz high-pass, 60 Hz notch, no low-pass), and resampled all data to 500 Hz before segmenting into 30-second windows. Please refer to Appendix F.3 for more details.
Scaling. As DIVER-1 extends the scale of EFMs, we require methods to transfer optimal hyperparameters across different model sizes. In standard parameterizations, optimal hyperparameters are highly dependent on model width. Maximal Update Parametrization ($\mu P$) (muparamyang2022tensor), by carefully scaling initializations and learning rates, ensures consistent weight update magnitudes as model width increases. This enables $\mu\text{Transfer}$, where optimal hyperparameters from small models can be directly transferred to large models, precluding expensive hyperparameter tuning.
DIVER-1 was pretrained on either 128, 48, 32 NVIDIA A100 GPUs or 32, 24, 16 H200 GPUs, depending on the experimental configuration. We tested models varying in number of parameters, from 12.72M to 1.83B as in Table 1. Additional training details are provided in Appendix section B.4. We conducted comprehensive scaling law experiments across four dimensions: compute (training epochs 1-64), model size (varying width, depth, and patch size), dataset size (1-100% of available data), and subject diversity (2-16 subjects with fixed dataset size). Detailed experimental configurations and scaling behaviors for each dimension are provided in the Appendix Table 1.

### v2

@@ Abstract
Unifying the vast heterogeneity of brain signals into a single foundation model is a longstanding challenge in neuroscience. Yet, even as large-scale pretraining becomes feasible, the field lacks principled guidance on how to scale EFMs under realistic data and compute constraints. We present the first systematic scaling law analysis spanning both EEG and iEEG and uncover a distinct “data-constrained” characteristic: unlike in language modeling, performance in electrophysiology is dominated first by data scale, followed by training duration (epochs), with model parameter count playing a subordinate role under fixed compute budgets. This challenges the prevailing “bigger is better” heuristic derived from LLMs. Building on these insights, we introduce DIVER-1, a family of models trained on the largest and most diverse corpus to date—59.3k hours (54k EEG + 5.3k iEEG) across 1.6M channel-hours from 17.7k+ subjects—scaling up to 1.82B parameters. [^1] Code available at: https://anonymous.4open.science/r/DIVER-1 By prioritizing data diversity and training horizons over mere parameter expansion, DIVER-1 achieves state-of-the-art performance across established benchmarks. Our work provides both a powerful generalist model and actionable guidelines for efficient development of future neuro-AI systems.
@@ 1 Introduction
Brain electrophysiology—scalp electroencephalography (EEG) and intracranial EEG (iEEG)—provides direct, millisecond-scale access to neural dynamics, enabling the study of rapid neural processes inaccessible to slower neuroimaging modalities (Luck, 2014; Cohen, 2014). It supports applications spanning neuroscience, brain–computer interfaces, and clinical monitoring, leveraging EEG/iEEG as both a research instrument and a clinical-grade measurement(Wolpaw et al., 2002; Tatum et al., 2022). However, electrophysiology is highly heterogeneous: signals vary across subjects, sessions, and brain states, and the sensing geometry itself changes across montages, channel counts, and electrode types. This heterogeneity is a primary barrier to transferable representations, and much recent electrophysiology foundation model (EFM) work has focused on objectives and architectures that explicitly cope with it rather than assuming a fixed sensor layout (Wang et al., 2024c; Zhang et al., 2023; Chau et al., 2025).
A natural hypothesis is that scaling—in data, compute, and model capacity—can help absorb electrophysiological heterogeneity into a single reusable backbone. Indeed, a small but growing set of EFMs has begun to probe “scaling,” typically through coarse ablations over model capacity and/or pretraining corpus size, and reports additional downstream gains with increased scale (Jiang et al., 2024; Wang et al., 2024b; c; 2023). However, it remains difficult to translate these findings into actionable guidance: explored regimes are still limited, and studies do not conduct compute-matched sweeps across multiple scaling axes (especially training duration). Moreover, it is unclear whether these trends hold consistently across both EEG and iEEG. This stands in contrast to language modeling, where scaling laws provide concrete guidance for trading off model size and data under a fixed compute budget to achieve strong performance (Kaplan et al., 2020; Hoffmann et al., 2022).
This gap is critical as most EFMs are developed with limited compute, where the cost of turning the wrong scaling knob is high. Additionally, ephys introduces domain-specific knobs beyond model and dataset size. In particular, training duration (number of epochs) becomes a first-class knob because ephys corpora are far from internet-scale, making repeated passes over the same recordings largely unavoidable. For iEEG, subject diversity introduces yet another knob, reflecting a trade-off between collecting data from more subjects versus collecting more recordings per subject. As a result, “how to scale” in electrophysiology is not a single decision but a knob-setting problem: how to allocate a fixed compute budget across model size, training duration, and data scale and composition to maximize transfer.
@@
[…]
These experiments reveals an actionable set of scaling knob settings: (i) turn the data knob as far as possible by acquiring more pretraining recordings, and (ii) under fixed data and compute budgets, turn the training-duration knob up—prioritizing longer training of smaller models over briefly training very large models (e.g., a 51M-parameter model trained for 32 epochs outperforms a 1.8B-parameter model trained for 8 epochs while using less compute). Importantly, these knob settings need not be chosen heuristically: the data-constrained scaling framework (Muennighoff et al., 2023) quantitatively predicts the compute-efficient allocation across model size, and epochs given fixed dataset size and compute. Guided by these insights, we introduce DIVER-1, a family of EEG and iEEG EFMs that achieves state-of-the-art performance with an architecture shaped by neuroscientifically motivated inductive priors; together with our scaling analysis, DIVER-1 provides a quantitative, modality-spanning scaling playbook for future EFM development.
The contributions of this work are threefold:
  • A quantitative scaling playbook for EFMs: We provide actionable guidance for compute-limited Ephys pretraining: (i) increasing data scale yields the largest gains, and (ii) under fixed data and compute, allocating more budget to epochs can be more beneficial than pushing to $\geq$1B-parameter models. We further show these trade-offs are precisely captured by data-constrained scaling laws (Muennighoff et al., 2023), which generalize classical Kaplan/Chinchilla-style scaling to the multi-epoch Ephys regime.
[…]
  • DIVER-1 model family with flexible electrode embeddings: We introduce DIVER-1 and achieve state-of-the-art performance across EEG and iEEG benchmarks using ephys-tailored components that better embed diverse electrode sets and make the encoder flexible to varying channel layouts.
2 DIVER-1 Architecture
Motivated by the central challenge of electrophysiology—robust generalization under heterogeneous and variable electrode configurations—we design DIVER-1 (Deep Integration of Vast Electrophysiological Recordings) around a single goal: flexible, informative embeddings for diverse electrode sets. DIVER-1 uses an architecture custom-designed for EEG and iEEG that supports effective self-supervised pretraining via masked patch reconstruction (Figure 1). The architecture consists of four components: (1) patch encoding, (2) embedding enhancements to inject modality- and channel-aware inductive priors, (3) a stack of MOIRAI blocks (Woo et al., 2024) to process variable-channel inputs, and (4) multi-output projection for multi-domain reconstruction. During pretraining, we randomly mask 50% of input patches and train the model to reconstruct the missing content across multiple signal domains through the projection layer. The details of pretraining and architectural hyperparameters can be found in Appendix B.6.
2.1 Patch Encoding
@@
We convert the raw signal into patch representations and extracts temporal features. Given an input time series $\mathbf{Y}\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels (variables) and $T$ is the sequence length, we patchify $\mathbf{Y}$ into 1 or 0.1 second patches, each containing $P\in\{500,50\}$ timepoints (at 500Hz sampling rate). During pretraining, $50\%$ of the patches are randomly masked with zero, and the remaining patches are passed through a three layer patch-wise convolutional neural network (CNN) to extract temporal features. This results in a basic representation $\mathbf{Y}_{\text{CNN}}\in\mathbb{R}^{C\times N\times d_{\text{model}}}$, where $N=T/P$ is the number of patches and $d_{\text{model}}$ is the embedding dimension.
2.2 Embedding Enhancements
This section introduces three sub-modules developed for improved learning over heterogeneous signal and channel information.
Spectral embedding ($\mathbf{E}_{\text{spectral}}$).
@@ 2.2 Embedding Enhancements
$\mathbf{E}_{\text{spectral}}$ injects patch-wise frequency information by applying FFT to each patch and projecting the features to $d_{\text{model}}$ dimensions via a learnable linear layer.
Channel position and modality embeddings.
@@
Both spatial electrode locations and electrode modality are jointly embedded to handle heterogeneous recording setups. While EEG follows standardized montages, iEEG channels are implanted in specific brain regions based on individual clinical needs. To address this heterogeneity, we encode both channel position embeddings $\mathbf{E}_{\text{position}}$ using the 3D spatial coordinates (registered to MNI space) and modality embeddings $\mathbf{E}_{\text{modality}}$ to distinguish between different electrode types. We then define the combined position–modality embedding term as:
$\mathbf{E}_{\text{position}},\mathbf{E}_{\text{modality}}]=\big[\big(PE^{(d)}(d)\big)_{d\in\{x,y,z\}},\;\mathbf{E}_{\text{modality}}\big],$ (1)
where electrode coordinates $(x,y,z)$ are in mm and:
@@
for $j\in\{x,y,z\}$ and $d_{x}=d_{y}=d_{z}=d_{\text{model}}/4$, following PopT (Chau et al., 2025). If the coordinates x,y,z for a given channel are not available, the positional embedding for that channel is set to zero. The modality embedding $\mathbf{E}_{\text{modality}}$ is computed as $\mathbf{E}_{\text{modality}}^{\text{type}}+\mathbf{E}_{\text{modality}}^{\text{subtype}}$, where $\mathbf{E}_{\text{modality}}^{\text{type}}$ distinguishes between EEG and iEEG channels, and $\mathbf{E}_{\text{modality}}^{\text{subtype}}$ encodes electrode subtypes (grid, strip, or depth for iEEG electrodes) through learnable embedding vectors shared across all patches within the same category.
Spatio-Temporal Conditional Positional Embedding (STCPE).
STCPE addresses the need for positional encodings that remain valid under heterogeneous and variable electrode configurations. Transformers require positional information because self-attention is inherently permutation-invariant. In vision, this is often handled with fixed positional encodings; in contrast, Conditional Positional Encoding (CPE) (Chu et al., 2021) injects position via lightweight convolutions over local neighborhoods, allowing the encoding to adapt to the structure of each input. Building on this idea, Asymmetric Conditional Positional Encoding (ACPE) (Wang et al., 2024c) applies asymmetric convolutions across channels and time for EEG. However, ACPE’s channel-wise convolutions are defined along a fixed channel axis and therefore do not preserve channel permutation equivariance: the learned spatial relationships depend on the training-time channel order. This is a critical limitation for electrophysiology, where robust transfer requires invariance to arbitrary electrode reorderings and flexibility across montages, recording systems, and channel counts.
@@
$\begin{split}\mathbf{E}_{\text{STCPE}}[:,t,:]&=P_{\uparrow}\!\left(\sum_{k=-m}^{m}\mathbf{H}_{t+k}[:,\,k+m,:]\right)\\ &\in\mathbb{R}^{C\times w\times(d_{\text{model}})},\qquad t=1,\dots,N.\end{split}$ (3)
@@
Overall inputs.
The final transformer input is computed as:
2.3 MOIRAI Block
Any-variate attention. Unlike vanilla attention mechanisms that rely solely on input embeddings to differentiate between tokens and model their relationships, any-variate attention (Woo et al., 2024) enables adaptive spatial modeling across heterogeneous electrode configurations by directly embedding spatio-temporal information into the attention computation itself through two main components: Rotary Position Embedding (RoPE) (Su et al., 2024) and binary attention bias.
@@ 2.3 MOIRAI Block
$\begin{split}E_{ij,mn}=(\mathbf{W}^{Q}\bm{x}_{i,m})^{T}\mathbf{R}_{i-j}(\mathbf{W}^{K}\bm{x}_{j,n})\\ +u^{1}\cdot\mathbbm{1}_{\{m=n\}}+u^{2}\cdot\mathbbm{1}_{\{m\neq n\}}\end{split}$ (5)
@@
Spatio-temporal register tokens. Following register tokens in vision transformers Darcet et al. (2023), we augment the $C\times N$ token grid with learnable registers for channel, temporal, and joint spatio-temporal information. This transforms input token grid shape from $C\times N\times d_{\text{model}}$ to $(C+1)\times(N+1)\times d_{\text{model}}$. These registers provide dedicated workspace for auxiliary computations without perturbing the primary Ephys representations.
2.4 Multi-Output Projection
We perform reconstruction across multiple signal domains during pretraining. Refined spatio-temporal representations are linearly projected to reconstruct patches in three complementary domains: raw time series, FFT, and STFT. This multi-domain reconstruction objective encourages comprehensive learning of temporal and spectral properties.
@@ 3.1 Pretraining
DIVER-1 employs self-supervised pretraining based on masked patch reconstruction.
Multi-domain reconstruction objective. Rather than reconstructing only raw time series, DIVER-1 utilizes a multi-output projection architecture that maps each learned patch representation $\mathbf{h}_{c,n}\in\mathbb{R}^{d_{\text{model}}}$ to three complementary signal domains through parallel linear transformations. For each patch $(c,n)$, the representation is projected to: (1) raw time series $\hat{\mathbf{y}}_{c,n}^{\text{raw}}\in\mathbb{R}^{P}$ to reconstruct temporal dynamics, (2) FFT coefficients $\hat{\mathbf{y}}_{c,n}^{\text{FFT}}\in\mathbb{R}^{P/2+1}$ to capture frequency domain characteristics, and (3) STFT spectrogram $\hat{\mathbf{y}}_{c,n}^{\text{STFT}}\in\mathbb{R}^{F\times T_{s}}$ to model time-frequency relationships, with specific FFT and STFT parameters detailed in Appendix B.5.
@@
$\mathcal{L}_{\text{total}}=\sum_{(c,n)\in\mathcal{M}}\ \sum_{k\in\mathcal{K}}\lambda_{k}\,\mathcal{L}_{\text{MSE}}\!\left(\mathbf{y}^{(k)}_{c,n},\hat{\mathbf{y}}^{(k)}_{c,n}\right)$ (6)
where $\mathcal{K}=\{\mathrm{raw},\mathrm{FFT},\mathrm{STFT}\}$ denotes the set of signal domains, $\lambda_{k}$ is the corresponding loss weight for domain $k$ and $\mathcal{M}$ denotes the set of masked patch indices. We used $(\lambda_{k})_{k\in\mathcal{K}}=(1,0.1,1)$ for 1 s patches model and $(1,1,0)$ for 0.1 s patches.
@@
Pretraining dataset. DIVER was pretrained on, to our knowledge, the largest and most diverse collection of Ephys datasets compared to previous EFMs. Our pretraining data encompasses diverse recording conditions including task-based experiments, resting-state recordings, and sleep studies, ensuring robust representation learning across different brain states. The pretraining datasets are summarized in Table 3 with further details in Appendix F.1. All data underwent QAQC, minimal preprocessing, and resampling with the goal of preserving as much of the original signal as possible. Please refer to Appendix F.3 for more details.
Scaling. As DIVER-1 extends the scale of EFMs, we require methods to transfer optimal hyperparameters across different model sizes. In standard parameterizations, optimal hyperparameters are highly dependent on model width. Maximal Update Parametrization ($\mu P$) (Yang et al., 2022), by carefully scaling initializations and learning rates, ensures consistent weight update magnitudes as model width increases. This enables $\mu\text{Transfer}$, where optimal hyperparameters from small models can be directly transferred to large models, precluding expensive hyperparameter tuning.
Training details. DIVER-1 was pretrained on either 128, 48, 32 NVIDIA A100 GPUs or 32, 24, 16 H200 GPUs, depending on the experimental configuration. We tested models varying in number of parameters, from 12.72M to 1.83B as in Table 2. Additional training details are provided in Appendix section B.4. We conducted comprehensive scaling law experiments across four dimensions: compute (training epochs 1-64), model size (varying width), dataset size (1-100% of available data), and subject diversity (2-16 subjects with fixed dataset size). Detailed experimental settings and scaling behaviors are provided in the Appendix.